Determine the coordinates of the centroids with respect to x and y axes. - 30 []] 15 12 065 Y 10 2 85
Answer #1

Solution 15 mm 21 30 mm 12mm 3 40mm so 10 mm X 85 mm By parts method & Αια Ě Ai 0 = EA; Y; E AI 201.. coordinate X - centroidX = 85 2 42.5 mm i x = 12 6mm x3 = 30 2 15mm 30mm Y = 0 = 5m 2 = 10 + 40 y3 = 10+ 40 + 15 2 = 57.5mm 2 X = A,a, & A 22 & Azaz